OpenStudy (anonymous):

Solve. Assume the exercise describes a linear relationship. When making a telephone call using a calling card, a call lasting 3 minutes costs $1.05. A call lasting 12 minutes costs $2.85. Let y be the cost of making a call lasting x minutes using a calling card. Write a linear equation that models the cost of making a call lasting x minutes.

OpenStudy (anonymous):

2.85 - 1.05/12 - 3 1.8 / 9 m = .2 y - 1.05 = .2(x - 3) y = .2x +.45